Justin is 57 years younger than Alma. 8 years ago, Alma's age was 4 times Justin's age. How old is Justin now?
Answer by ankor@dixie-net.com(22740)

Justin is 57 years younger than Alma.
j = a - 57
or
a = j + 57
8 years ago, Alma's age was 4 times Justin's age.
a - 8 = 4(j-8)
a - 8 = 4j - 32
a = 4j - 32 + 8
a = 4j - 24
How old is Justin now?
replace a with j + 57
j + 57 = 4j - 24
57 + 24 = 4j - j
81 = 3j
j = 81/3
j = 27 yrs is J's age now
:
:
Check this, find A's age and check in the 2nd statement
27 + 57 = 84
84 - 8 = 4(27-8)
76 = 4 * 19